### v2

# Understanding thermalization in a non-Abelian gauge theory in terms of its soft modes

Abstract
We measure the maximal Lyapunov exponent $\lambda_{L}$ of physical states in a SU(2) gauge theory consisting of soft momentum modes both in and out-of-thermal equilibrium conditions using ab-initio lattice techniques. We have implemented different algorithms to appropriately describe the dynamics of soft-modes for a wide range of temperatures and under non-equilibrium conditions. The non-equilibrium state has been realized starting from an over-occupied initial condition for low momentum soft gluons whereas the thermal state comprises of strongly interacting soft gluons at temperatures where these are well separated from the hard momentum modes. Spectra of positive Lyapunov exponents is observed in both these states, similar to a chaotic dynamical system. From the Kolmogorov-Sinai entropy rate measured in terms of this spectrum, we estimate a typical time-scale of $\sim 0.50(3)$ fm/c to achieve thermalization at $T\sim 600$ MeV starting from the non-thermal state. We also measure, for the first time, the $\lambda_{L}$ for long wavelength critical modes of SU(2) using the out-of-time-ordered correlator of a classical $Z_{2}$ scalar field theory, which shares the same universal behavior with SU(2), near the deconfinement phase transition. The $\lambda_{L}$ is observed to maximize at the transition temperature.
keywords:
Non-equilibrium quantum field theory , QCD , non-Abelian gauge theory , chaos , out-of-time-ordered correlator
journal: Physics Letters B
[…]
[first]organization=The Institute of Mathematical Sciences, a CI of Homi Bhabha National Institute,city=Chennai, postcode=60113, country=India
1 Motivation
Understanding the mechanism of thermalization in strongly interacting systems is a challenging research problem in theoretical physics [1, 2] as it would be able to explain physics at different length scales, e.g. in heavy-ion collisions, cold atom traps or in the early universe. Starting from a non-equilibrium state, a necessary condition for eventually reaching a thermal state is a mixing between different regions of the phase space. A mixing in the phase space can occur if there is a mechanical instability in the system [3], which is characterized by at least one positive Lyapunov exponent. Furthermore in the thermal state the entropy or the information content [4] is maximized [5]. Except for a few simple systems [6, 7] it is not very well understood how the entropy evolves as a function of time, starting from a chaotic non-equilibrium state and eventually maximizing itself in a thermal state. One important observable is the rate of change of entropy, known as the Kolmogorov-Sinai (KS) entropy rate [8], which equals the sum of all positive Lyapunov exponents for Hamiltonian systems [9]. This equality is known to be true for dissipative systems that have an attractor solution. Hence calculating the Lyapunov exponents in chaotic dynamical systems is crucial to understand if, and in what time-scales would these eventually thermalize.
In this Letter, we revisit the question whether non-Abelian gauge theories are chaotic dynamical systems [10, 11, 12], however, with a goal to understand the phase space mixing and the process of entropy maximization in an interacting quantum field theory. For this purpose, we measure the maximal Lyapunov exponent of 4-dimensional SU(2) gauge theory using ab-initio lattice techniques in three distinct regimes; for a particular non-thermal classical attractor state [13, 14] and for thermal states near to the color-deconfinement phase transition and at high temperatures, deep in the deconfined phase. Measuring the KS entropy rate of the soft-gluon modes in a typical thermal state at $\sim 600$ MeV from the Lyapunov exponents, we for the first time provide an estimate of the typical thermalization time $\sim 0.50(3)$ fm/c for the system to relax into it, starting from the classical attractor regime. This estimate is consistent with the value that can be obtained using a recently proposed procedure of matching the scales that characterizes the infrared magnetic modes of QCD at equilibrium and in a non-thermal state in the self-similar scaling regime [15]. Apart from reinforcing the stochastic features of non-Abelian gauge theories [16] already strongly developed in its classical modes, our study has a practical relevance as well. Measuring the thermalization time of the chaotic soft-gluon modes thus provides an estimate of how long it takes to achieve local thermal equilibration i.e. to evolve to a quark gluon plasma phase at temperatures $450$-$600$ MeV, a range which is similar to the experimentally measured initial temperature of the fireball formed in a typical heavy-ion collision event [17, 18], starting from a highly occupied gluon phase.
2 Theoretical background
We begin with a discussion of how to measure the (maximal) Lyapunov exponent in a non-Abelian SU(2) gauge theory which is a non-linear quantum field theory. We recall that, in a classical dynamical system, starting from two closely spaced initial coordinates and performing a Hamiltonian evolution, if the separation between the two trajectories increases exponentially with time then the dynamics is defined as chaotic. The rate at which the trajectories diverge from each other is characterized by the Lyapunov exponent, $\lambda_{L}=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\frac{d(t)}{d(0)}$ where $d(t)$ denotes the separation between two classical trajectories at time $t$. This idea can be easily generalized in a quantum field theory like SU(2), in the semi-classical regime, where the gauge links $U_{\mu}(\mathbf{x})$ are analogous to the space coordinates in a classical configuration space. Using non-perturbative lattice techniques one can generate ensembles of gauge links using Monte-Carlo algorithms. In thermal equilibrium, two different gauge configurations which are separated by infinitesimal Monte-Carlo time steps can be evolved in real-time using the SU(2) Hamiltonian in order to estimate the Lyapunov exponent.
Alternately, another definition of the Lyapunov exponent can also be motivated by studying the following out-of-time ordered correlator (OTOC) defined in one-dimensional classical dynamics as,
$\{x(t),p(0)\}^{2}=\left(\frac{\delta x(t)}{\delta x(0)}\right)^{2}\sim\rm{e}^{2\lambda_{L}t}~{},$ (1)
The OTOCs can be easily generalized for quantum field theories. For real-valued scalar fields, the expression in the LHS of Eq. 1 can be re-written by replacing the Poisson bracket in terms of position and momenta with a commutator of the field operator $\hat{\phi}(\mathbf{x},t)$ and their conjugate momentum fields $\hat{\pi}(\mathbf{x},t)$. However for scalar fields the OTOC can also be equivalently defined as $[\hat{\phi}(\mathbf{x},t),\hat{\phi}(\mathbf{0},0)]^{2}$ [19]. In a quantum system, OTOC grows exponentially upto some time $t_{E}$, called the Ehrenfest time [20] beyond which its growth saturates. This is the timescale during which the wavefunction has spread over the entire volume of the quantum system and there is a transition from a wave-like to a particle-like behavior [21]. However in the semi-classical limit, the OTOC will not saturate. Such a scenario is realized when scalar fields have large occupation numbers, where one can replace the commutator with the following expression,
$C(\mathbf{x},t)=\langle\{\phi(\mathbf{x},t),\phi(\mathbf{0},0)\}^{2}\rangle=\langle\left(\frac{\delta\phi(\mathbf{x},t)}{\delta\pi(\mathbf{0},0)}\right)^{2}\rangle~{}.$ (2)
We henceforth denote the classical fields $\phi(\mathbf{x},t)$ as $\phi_{\mathbf{x}}(t)$ and similarly for its conjugate fields. If indeed the system is chaotic, the (maximal) Lyapunov exponent $\lambda_{L}$ can be extracted from the OTOC using the relation,
$\lambda_{L}=\lim_{t\rightarrow\infty}\frac{1}{2t}\ln~{}C(\mathbf{x=0},t)~{}.$ (3)
An OTOC typically spreads ballistically within the light cone along rays of constant speed $\mathrm{v}=x/t$ in the $x$-$t$ plane with a scaling exponent $\nu$ [22] as,
$C(\mathbf{x},t)\sim\rm{e}^{2\lambda(v)t}~{},~{}\lambda(v)=\lambda_{L}~{}\left[1-\left(\frac{v}{v_{B}}\right)^{\nu}\right]~{}.$ (4)
From the spectrum of the Lyapunov exponents $\lambda(\mathrm{v})$ in the velocity space of propagating OTOCs, one can define the butterfly velocity $\mathrm{v}_{B}$ from the relation $\lambda(\mathrm{v}_{B})=0$. The quantities $\lambda_{L},\mathrm{v}_{B}$ and exponent $\nu$ are all, in general, temperature dependent. The diffusion coefficient $D$ is related to the maximal Lyapunov exponent $\lambda_{L}$ and the butterfly velocity $\mathrm{v}_{B}$ through the relation [23],
$D=\frac{\mathrm{v}_{B}^{2}}{\lambda_{L}}~{}.$ (5)
In Section 4, we will extract $\lambda_{L},\mathrm{v}_{B}$ and $D$ for SU(2) gauge theory in thermal equilibrium for a wide range of temperatures as well as in a particular non-equilibrium state. The ab-initio lattice algorithms used for this purpose are detailed in the next section.
3 The algorithm
3.1 Studying highly occupied SU(2) gauge theory under non-equilibrium conditions
For a specific realization of a non-equilibrium state of SU(2) gauge theory, we start from an initial condition where modes of gauge fields the momentum $\mathbf{p}$ are occupied according to a phase-space distribution given by, $\tilde{f}(\mathbf{p})=g^{2}f(\mathbf{p})=n_{0}\frac{Q_{s}}{|\mathbf{p}|}\rm{e}^{\frac{-|\mathbf{p}|^{2}}{2Q_{s}^{2}}}$  [15]. Here $Q_{s}$ is the gluon-saturation scale which is typically between $1$-$2$ GeV and $g$ is the gauge coupling. We consider $Q_{s}=1.5$ GeV in this work. The choice of the initial phase-space distribution becomes irrelevant after a time scale $Q_{s}t\sim n_{0}^{-2}$, which corresponds parametrically to the inverse of the large angle scattering rate between the quasiparticles [24, 25, 14]. Here and in the subsequent sections we would be considering the theory to be discretized on a 3-dimensional spatial box with $N$ sites along each spatial direction and $a$ as the lattice spacing. The quantity $n_{0}$ sets the initial occupation number of the infrared modes of the gauge fields which we have chosen such that it is close to the typical equilibrium energy densities at high temperatures $T\gtrsim 600$ MeV. Since the occupation numbers are large (owing to g being small), the gauge links behave classically. For generating these classical configurations, we solve for the Hamilton’s equations of motion for the gauge links $U^{i}_{\mathbf{x}}(t)$ and its conjugate fields $E^{ib}_{\mathbf{x}}(t)$ on the lattice along directions $i=1,2,3$ and with color components $b=1,2$,
$E^{ib}_{\mathbf{x}}\left(t+\frac{\Delta t}{2}\right)$ $=$ $E^{ib}_{\mathbf{x}}\left(t-\frac{\Delta t}{2}\right)+2\frac{\Delta t}{a}\sum_{j\neq i}^{N^{3}}\text{tr}\left[i\Gamma^{b}(V^{ij}_{\mathbf{x}}+c.c)\right]~{},$
$U^{i}_{\mathbf{x}}(t+\Delta t)$ $=$ $\rm{e}^{i\frac{\Delta t}{a}~{}E^{ib}_{\mathbf{x}}\left(t+\frac{\Delta t}{2}\right)\Gamma^{b}}U^{i}_{\mathbf{x}}(t)~{},$ (6)
where $\Gamma^{b}$ are generators of SU(2) and $V^{ij}_{\mathbf{x}}$ is the spatial staple defined in the $i$-$j$ plane as the derivative of the gauge plaquette $U_{P}$, defined using path ordering $\mathcal{P}$ in color space as,
$V^{ij}_{\mathbf{x}}=\frac{\delta U_{P}}{\delta U^{i}_{\mathbf{x}}}~{},~{}U_{P}^{ij}=\mathcal{P}\left(U^{i}_{\mathbf{x}}U^{j}_{\mathbf{x+\hat{i}}}U^{i{\dagger}}_{\mathbf{x+\hat{j}}}U^{j{\dagger}}_{\mathbf{x}}\right).$ (7)
The initial conditions for the gauge and electric fields are chosen as a classical-statistical ensemble representing vacuum fluctuations of transversely polarized gluons [24]. We use a symplectic leap-frog integrator for the time evolution in order to respect time-reversal invariance, with a time-step $\Delta t=0.0125Q_{s}$. We study different lattice volumes with extent along the spatial directions $N=128$, with spacings $aQ_{s}=1.115,0.9$ such that corresponding effective temperatures match with the thermal states at temperature $T$ which were generated using the algorithm discussed in the next section i.e. $T_{\mathrm{eff}}\approx T$. The spatial length of the box in physical units is $\sim 16$ fm. The Gauss law $\frac{a}{\Delta t}\sum_{i}\text{tr}i\Gamma^{a}\left[U_{\mathbf{x}}^{i}(t)-U^{i}_{\mathbf{x}+\hat{i}}(t)\right]$ was implemented with a precession of $10^{-15}$ at $t=0$ and it was checked to remain so at later times. With such a choice of initial condition the system eventually enters a self-similar scaling regime at late times where the gluon distribution function takes the form $\tilde{f}(\mathbf{p},t)=(Q_{s}t)^{-4/7}f_{s}((Q_{s}t)^{-1/7}\mathbf{p})$, where $f_{s}$ is a homogeneous function describing a static distribution [14]. Within such a regime the hard, electric and the magnetic momentum scales, the latter two corresponding to scales set by the Debye mass and the square root of the spatial string tension respectively, are well separated similar to a thermal plasma [26].
In order to measure the degree of chaoticity in such a system in a gauge-invariant manner, we follow the procedure as outlined in Ref. [15]. First we implement an infinitesimal shift in the initial condition to $n_{0}+\delta n_{0}$ where $\delta n_{0}=10^{-5}$ and perform a Hamiltonian evolution of the gauge links $U^{\prime}_{\mathbf{x}}(t)$ as a function of time. We next measure the separation between two gauge trajectories at a time $t$ starting with these two infinitesimally close initial conditions, through the gauge-invariant distance measure defined in terms of the spatial plaquettes,
$d(t)=\frac{1}{2N_{p}}\sum_{P}|\text{tr}U_{P}(t)-\text{tr}U^{\prime}_{P}(t)|\sim e^{\lambda_{L}t}$ (8)
where $N_{P}$ are the number of independent plaquettes. We particularly choose the time interval to be within the self-similar scaling regime, in order to extract the Lyapunov exponent.
3.2 Studying thermal SU(2) plasma at high temperatures
At finite temperatures, apart from the hard scale $\pi T$ present naturally in the problem, there are two new scales that are generated dynamically: the electric scale $gT$ and the magnetic scale $g^{2}T/\pi$ [27]. When these three scales are well-separated, one can integrate out the hard and electric gluons with momenta $\gtrsim gT$ to obtain an effective Hamiltonian [28], which describes the dynamics of the magnetic (soft) gluons with the color electric fields $\tilde{E}_{\mathbf{x}}^{i}$ satisfying the equation of motion
$-\partial_{t}\tilde{E}_{\mathbf{x}}^{i}+[D_{j},\tilde{F}^{ji}(\mathbf{x})]=\sigma\tilde{E}^{i}_{\mathbf{x}}+\tilde{\zeta}^{i}_{\mathbf{x}}(t)~{},$ (9)
where $\sigma$ is the color conductivity and $\tilde{\zeta}^{i}_{\mathbf{x}}(t)$ are the stochastic color-force fields which satisfy $\langle\tilde{\zeta}^{ic}_{\mathbf{x_{1}}}(t_{1})\tilde{\zeta}^{jb}_{\mathbf{x_{2}}}(t_{2})\rangle=2T\sigma\delta^{ij}\delta^{cb}\delta^{3}(\mathbf{x_{1}-x_{2}})\delta(t_{1}-t_{2})$ in accordance to the fluctuation-dissipation relation. Discretizing the Eq. 9, the electric fields, noise fields and color-conductivity can be written in dimensionless units on the lattice in terms of $E_{\mathbf{x}}^{i}=g\tilde{E}_{\mathbf{x}}^{i}a^{2},~{}\zeta^{i}_{\mathbf{x}}=a^{3}g\tilde{\zeta}^{i}_{\mathbf{x}}(t)$ and $\sigma a=\frac{\sigma}{T}.Ta$ respectively. We set $\sigma/T$ to its perturbative values [29], given by the relation
$\sigma^{-1}=\frac{3N_{c}g^{2}T}{4\pi m_{D}^{2}}\left[\ln\frac{m_{D}}{\gamma}+3.041\right]~{},$ (10)
where the coupling $g$ at each temperature $T$ is determined using the 2-loop $\beta$ function and $1/\gamma$ is the mean free path of the color changing processes which varies between $\gamma/m_{D}\in[0.883,1.034]$ for the range of temperatures considered in our work. Here $N_{c}=2$ and the Debye mass is defined at one-loop order in perturbation theory i.e., $m_{D}=\sqrt{\frac{2}{3}}gT$. We numerically implement Eq. 9 in dimensionless units which is similar to the electric field evolution in Eq. 6 but with additional contributions due to the noise and damping terms given by $-\Delta t/a\left[a\sigma E^{ib}_{\mathbf{x}}(t-\Delta t)-\mathbf{\zeta}^{ib}_{\mathbf{x}}(t-\Delta t)\right]$. The evolution in Eq. 9 is similar to a Langevin equation for magnetic modes of the gauge fields in presence of a heat bath. Here, the noise term represents the effects of the hard modes which form the heat bath and the damping dynamically brings the system to a thermal equilibrium. This classical system consists of $6N^{3}$ oscillators of energy $aT$ and has an energy density on the lattice [12] given by $\frac{6}{N^{3}}\sum_{\mathbf{k}}|\mathbf{k}|\frac{aT}{|\mathbf{k}|}~{},$ where the sum is over all allowed lattice momenta $\mathbf{k}$, which we have also verified in our calculations. On the other hand, energy density of the quantum SU(2) gauge theory at temperatures $\sim 2$ times the deconfinement temperature is close to its Stefan-Boltzmann limit $\pi^{2}T^{4}/5$ [30]. The lattice spacing in the effective theory is set in physical units using the criterion that the measured energy density matches with the quantum theory at each $T$. This results in a condition $T.a=(30/\pi^{2})^{1/3}$, which we use to set the lattice spacing in physical units. Once the gauge links are thermalized at a temperature $T$, there is an onset of a plateau in the values of the plaquette as a function of Langevin time. We then consider two thermal configurations infinitesimally separated as a function of Langevin time, switch off the noise and the damping terms in the Hamiltonian and evolve them subsequently as a function of real-time. This is to implement non-equilibrium evolution of the system of soft gluons along the Schwinger-Keldysh contour, which for this case, consists of only forward-in-time propagation since soft gluons have large occupation numbers. We then calculate the gauge invariant distance measure $d(t)$ introduced earlier to calculate the (maximal) Lyapunov exponent.
3.3 Chaotic dynamics near deconfinement phase transition
Near the deconfinement phase transition temperature $T_{c}$ the hard, electric and magnetic scales coincide, hence the effective thermal field theory described in the preceding section can no longer describe the gauge field dynamics. Hence we change our strategy to estimate the Lyapunov exponent. Since SU(2) gauge theory shares the same universality class as the 3D Ising model [31], hence its long-wavelength or low energy critical modes will undergo the same dynamics as a $Z_{2}$ scalar field. We thus simulate a relativistic $Z_{2}$ scalar field theory to understand the chaotic dynamics of the critical modes near the deconfinement phase transition. We start with the Hamiltonian defined on the lattice as,
$H=\sum_{\mathbf{x}}\left[\frac{\pi_{\mathbf{x}}^{2}}{2}-\frac{1}{2}\sum_{\mathbf{y\sim x}}\phi_{\mathbf{x}}\phi_{\mathbf{y}}+\left(\frac{m^{2}a^{2}}{2}+3\right)\phi_{\mathbf{x}}^{2}+\frac{\lambda_{s}}{4!}\phi_{\mathbf{x}}^{4}\right]~{}.$ (11)
Here $\sum_{\mathbf{y\sim x}}$ implies the sum over all nearest neighbors $\mathbf{y}$ around the lattice site $\mathbf{x}$. In our work, we set the values of $m^{2}a^{2}=-1~{},~{}\lambda_{s}=1$. We first generate thermal ensembles using the Langevin dynamics, in which we consider the time evolution of the scalar and its conjugate momentum fields through equations of motion,
$\partial_{t}\phi_{\mathbf{x}}(t)$ $=$ $\frac{\partial H}{\partial\pi_{\mathbf{x}}}~{}=~{}\pi_{\mathbf{x}}(t)~{},$ (12)
$\partial_{t}\pi_{\mathbf{x}}(t)$ $=$ $-\frac{\partial H}{\partial\phi_{\mathbf{x}}}-\gamma\pi_{\mathbf{x}}(t)+\eta_{\mathbf{x}}(t)~{}\sqrt{2\gamma T}~{}.$ (13)
where $\gamma$ is the damping term and $\eta_{\mathbf{x}}$ are Gaussian noise fields defined on each lattice site with a zero mean value and satisfying $\langle\eta_{\mathbf{x}}(t)\eta_{\mathbf{y}}(t^{\prime})\rangle=\delta_{\mathbf{xy}}\delta(t-t^{\prime})$. The relation between the damping factor $\gamma$ and the coefficient of the noise term mimicking thermal fluctuations can be obtained using the fluctuation-dissipation theorem. To initialize the algorithm, we select random initial values of $\phi_{\mathbf{x}}(0)$ and $\pi_{\mathbf{x}}(0)$ on each lattice site $\mathbf{x}$. We then numerically evolve the above differential equations using the symplectic leap-frog integrator with a time step $\Delta t=0.01$. Since Langevin dynamics is ergodic and follows the condition of detailed balance, we expect the system to thermalize after some Langevin time $t_{L}$. Once we achieve thermalization using the Langevin dynamics, we take the thermal configurations of the scalar fields and perform a real-time classical evolution. Further, we implement infinitesimal perturbation about the initial thermal configurations, denoted by $\delta\phi_{\mathbf{x}}(0)=0$ and $\delta\pi_{\mathbf{x}}(0)=r\delta_{\mathbf{x0}}$, where $r$ is a Gaussian random number between 00 and $0.01$. Since $\delta\pi_{\mathbf{x}}(t)=\partial_{t}\delta\phi_{\mathbf{x}}(t)$ one can re-express the initial momentum profile in the configuration space at $\mathbf{x}\equiv(0,0,0)$ and $t=0$ and then evolve the fluctuations of the scalar field via
$\partial_{t}^{2}\delta\phi_{\mathbf{x}}(t)=\nabla^{2}\delta\phi_{\mathbf{x}}(t)-m^{2}\delta\phi_{\mathbf{x}}(t)-\frac{\lambda_{s}}{2}\phi^{2}_{\mathbf{x}}(t)~{}\delta\phi_{\mathbf{x}}(t)~{}.$ (14)
In Eq. 14 one requires the values of the scalar fields $\phi_{\mathbf{x}}(t)$ as an input which are obtained through a classical Hamiltonian evolution of initial thermalized values of $\phi_{\mathbf{x}}(0)$ at each temperature $T$. We further evolve the initial momentum field fluctuations $\delta\pi_{\mathbf{x}}(t)$ using $\partial_{t}\delta\phi_{\mathbf{x}}(t)=\delta\pi_{\mathbf{x}}(t)$. We thus have all the ingredients to calculate OTOC in the semiclassical limit using Eq. 2, from which we extract the Lyapunov exponent, as discussed in the next section.
Figure 1: The Lyapunov exponent as a function of $T/T_{c}$ for a $Z_{2}$ scalar field theory, which from universality represent the critical long-distance modes of a $SU(2)$ thermal state near the deconfinement transition temperature $T_{c}$. The inset shows the growth of OTOC with time $t.T_{c}$.
4 Results
We first measure the OTOCs for $Z_{2}$ scalar field theory at different temperatures close to the phase transition, shown in the inset of Fig 1. After an initial dip, the OTOCs exhibit an exponential rise as a function of time. By performing a fit to the OTOCs at late times $t.T_{c}>1000$ with an exponential ansatz, we extract the (maximal) Lyapunov exponent at different temperatures, results of which are shown in Figure 1. The Lyapunov exponent thus extracted varies as $\lambda_{L}\sim T^{6.8(7)}$ below $T_{c}$, whereas in the range $(1,1.05)T_{c}$ its temperature variation can as characterized as $T^{-2.8(3)}$, which was determined after performing a fit to our data. In our calculations $a.T_{c}=9.37$ which agrees very well with the values obtained in the literature [32]. For reference, the deconfinement phase transition temperature in SU(2) is $T_{c}\sim 300$ MeV [30].
Further, in the inset of Fig 2, we show the heat-map of the OTOC at $T=T_{c}$ in a 2-dimensional plane as a function of time $t$ and one of the spatial coordinates, performing a sum over the other spatial coordinates. The OTOC spreads ballistically within the light cone from an initial small perturbation at the origin as evident from the color gradient of the heat-map. In general, $\lambda_{L}$ is a function of the velocity and it can be measured from the OTOC using Eq. 4. As evident in Fig 1, the Lyapunov exponent is maximum along $\mathrm{v}=0$ i.e. along the time direction where the rate of increase in the velocity is the largest compared to any other direction. We next calculate the butterfly velocity $\mathrm{v}_{B}$ by fitting the OTOC $C(\mathbf{x},t)$ to the ansatz given in Eq. 4. The $\mathrm{v}_{B}$ varies across the phase transition and is found to be maximum at $T=T_{c}$. In Fig 2, we have shown the diffusion coefficient as a function of temperature, which is related to the butterfly velocity through the relation in Eq. 5. As evident from the figure, the diffusion coefficient is constant below $T_{c}$ but decreases with increasing temperatures as $(T/T_{c})^{-3.7(2)}$ for $T>T_{c}$.
Figure 2: The temperature dependence of the diffusion coefficient $D$ for $Z_{2}$ scalar fields near $T_{c}$, extracted using Eq. 4. The inset shows the ballistic spread of the OTOC of $Z_{2}$ scalar fields in $x$-$t$ plane at $T_{c}$. The color profile from green to red denotes increasing values of OTOC in the $x$-$t$ plane.
At higher temperatures, $T\gg T_{c}$, the gauge invariant distance measure $d(t)$ was measured on a lattice with $N=32$ spatial sites and for different temperatures, results for which are shown in the inset of Fig 3. The initial $d(t)\sim 10^{-15}$ measures the typical difference between the plaquettes in two thermal configurations which, when evolved in time, grows exponentially eventually saturating at late times to $\sim 10^{-2}$, since the trace of plaquette is bounded to unity and the gauge group is compact. Within this timescale, the momentum mode occupancies of gluons remain similarly close to the thermal distribution. The lattice size is chosen to be large enough such that the physical lengths are $\sim 15(6)$ fm at temperatures $0.6(1.5)$ GeV respectively. We extract the (maximal) Lyapunov exponent $\lambda_{L}$ from the exponentially rising part of $d(t)$ and study its temperature dependence, results of which are shown in Fig. 3. The $\lambda_{L}$ varies linearly with temperature in the range $0.6$-$3$ GeV with a slope $\lambda_{L}/T\sim 0.52$ which satisfies the conjectured bound $\lambda_{L}\leq 2\pi T$ in quantum many-body chaotic systems [33] and also satisfies $\lambda_{L}\sim g^{2}\mathcal{E}/6$, where $\mathcal{E}$ is the average energy per plaquette [11]. The diffusion coefficient extracted from $\lambda_{L}$ using Eq. 5 varies inversely with temperature as $(T/T_{c})^{-0.95(3)}$ shown in Fig  4. At these temperatures the distance measure spreads from an initial tiny fluctuation at $t.T=0$ spreads ballistically over the space-time and grows exponentially within the light-cone, evident from the inset of the same figure, the color gradient from green to red indicates increasing values of the distance measure $d(t)$. Across these high temperatures, the butterfly velocity is $\mathrm{v}_{B}=0.8c$ and does not vary significantly. Compared to its value at $T_{c}$, the butterfly velocity reduces by only $\sim 10\%$ which implies that once thermalized, most of the configuration (color) space gets occupied.
Figure 3: The Lyapunov exponent as a function of $T/T_{c}$ for a SU(2) gauge theory in thermal equilibrium at high temperatures $T>>T_{c}$, shown as triangles. These are compared with the Lyapunov exponents in a non-thermal state of SU(2) with comparable energy densities (circles). The inset shows the growth of the gauge invariant distance measure with time $t.T_{c}$ for a thermal state of SU(2) for 3 different temperatures.
We next extract the Lyapunov exponent within the self-similar scaling regime of the non-equilibrium SU(2) plasma as outlined in Sec. 3.1. In order to compare with a typical thermal state at temperature $T\sim 2$ GeV, we set the initial density of gluons to be $n_{0}=16$, and consider lattice box of spatial extent $\sim 16$ fm such that the energy densities are similar in the thermal as well in the non-thermal scaling regime. The (maximal) Lyapunov exponent $\lambda_{L}=0.66~{}Q_{s}$ extracted in this non-thermal state is similar in magnitude to that in a thermal state at $T\sim 2$ GeV. We have also shown the values of the maximal Lyapunov exponents in Fig 3 in this non-thermal state, as a function of the effective temperature extracted from the fourth root of the energy density, as blue data points. The data points lie on the same curve that characterizes the temperature dependence of the thermal Lyapunov exponents. For lower energy densities $\varepsilon(\sim\sqrt{n_{0}}\lesssim 1)$ it was earlier conjectured [34] and later observed [15] from classical-statistical lattice calculations in the self-similar non-thermal scaling regime of SU(3), that the (maximal) Lyapunov exponent $\lambda_{L}$ varies as $\varepsilon^{1/4}$. Our calculations performed in the non-thermal scaling regime of SU(2), with a larger initial gluon energy density $(\sim\sqrt{n_{0}}\gtrsim 4)$ also yield a linear scaling of $\lambda_{L}$ with $\varepsilon^{1/4}$. The linear scaling of $\lambda_{L}$ at higher energies similar to a thermal plasma can be understood in terms of linear temperature dependence of the plasmon damping rate [35].
5 Physical implications of our results
Our results demonstrate that the infrared (soft) modes of SU(2) gauge theory exhibit chaotic dynamics both in thermal equilibrium at a wide range of temperatures $0.9$-$10~{}T_{c}$ as well in a particular non-thermal fixed point as evident from the positive (maximal) Lyapunov exponent $\lambda_{L}$. Interestingly the $\lambda_{L}$ decreases beyond the critical temperature $T_{c}$ with temperatures $T>T_{c}$, maximizing at the deconfinement phase transition. Such a property of $\lambda_{L}$ for the critical modes has been observed earlier in spin systems  [36, 37, 38, 39] as well as in scalar field theories in 2 dimensions [19]. The temperature dependence of $\lambda_{L}$, thus, is not particularly related to the details of microscopic interactions between the degrees of freedom but to the symmetries of the system.
Figure 4: The temperature dependence of the diffusion coefficient $D$ in SU(2) gauge theory at high temperatures $T\gg T_{c}$. The inset shows the ballistic spread of the Lyapunov exponents in a typical thermal state of SU(2) at $T=1.2$ GeV. The color gradient from green to red denotes increasing values of the distance function in the $x$-$t$ plane.
The diffusion coefficient $D$ which measures how fast the correlation between the fields increases with time, has many interesting features as a function of temperature. The $\lambda_{L}$ at $T_{c}$ shown in Fig. 1 correspond to the long-wavelength critical modes of SU(2) gauge theory which are a subset of the infrared magnetic modes, whose momenta are bounded by $|\mathbf{p}|\leq g^{2}T/\pi$. The fact that $D$ is independent of temperature below and at $T_{c}$ and decreases sharply above $T_{c}$ implies that these critical gluon modes at the phase transition are maximally spread in the configuration space. This spread is an inherently chaotic phenomena, which also manifests itself in the stochastic properties [40] of the eigenvalue density of Wilson loop operator in the limit of large number of colors. In this limit, the quantum fluctuations are sub-dominant, and an order-disorder transition [41, 42] in terms of the area of Wilson loops has similar features as the (de)confinement transition.
The temperature dependence of $D$ at $T\gg T_{c}$ as shown in Fig. 4 can be interpreted from the fact that $D$ represents a diffusion in the configuration space of color gauge links and hence can be related to the inverse of color conductivity $\sigma$. The color conductivity is related to the color diffusion coefficient through the relation, $\sigma=\chi D$ [43, 44], where $\chi$ is the color charge susceptibility. The $\sigma$ calculated using perturbative techniques in a thermal non-Abelian plasma, increases linearly with temperature [29] and $\chi$ varies as $\sim T^{2}$. Hence the diffusion coefficient varies as $D\sim\frac{1}{T}$, which parametrically depends on the inverse of color conductivity $\sigma$. In the effective dynamics of the soft (magnetic) modes described by Eq. 9, $\sigma$ represent how these modes rearrange in response to the ultra-relativistic hard modes. The soft modes receive random kicks from the thermal bath consisting of hard modes and a finite $\sigma$ effectively dampens them, allowing for efficient thermalization.
Next we address the implications of our study for understanding the mechanism of thermalization in gauge theories. When we discuss about thermalization, to be more precise, we imply how fast the magnetic (soft) modes equilibrate. Starting from a non-equilibrium state the entropy increases until it reaches its maximum in a thermal state. The KS entropy density rate $\dot{s}_{KS}$ can thus be used to estimate the time required in this process. From the sum of all the positive Lyapunov exponents $\lambda_{i}$ [9], one can estimate this quantity on the lattice as
$\dot{s}_{KS}a^{4}=-\frac{\sum_{i}\lambda_{i}a}{N^{3}}\sim-4\lambda_{L}.{a}~{},$ (15)
for the range of temperatures $2$-$10~{}T_{c}$, considering that about $1/3$ of $18N^{3}$ number of the Lyapunov exponents are positive [12] and their values can be obtained from the relation $\lambda(v)\sim\lambda_{L}[1-(\mathrm{v}/\mathrm{v}_{B})^{2}]$ evident in our data, where we have taken the values of $\mathrm{v}$ as randomly distributed between 00 and $\mathrm{v}_{B}$. We have also verified that our estimate does not vary significantly if the values of $\mathrm{v}$ are chosen to be uniformly distributed between $[0,\mathrm{v}_{B}]$. Starting from a thermal state, $\dot{s}_{KS}$ measures the rate at which information about the initial state of the system is lost. We use the thermodynamic relation $s_{\text{th}}T=\varepsilon+p$ to obtain the entropy density of a thermal state of SU(2) to be $s_{\text{th}}a^{3}=\frac{4}{3}\frac{\varepsilon.a^{4}}{T.a}$ at temperatures $T\gg T_{c}$ where the energy density and pressure are related by $\varepsilon\simeq 3p$. On the other hand the entropy density for the non-thermal state in the self-similar regime described in section 3.1, $s_{\text{non-th}}a^{3}=-\int{\tilde{f}(|\mathbf{p}|)\ln\left[\tilde{f}(|\mathbf{p}|)\right]~{}d^{3}(a\mathbf{p})}\simeq 4.8$ can be calculated by performing a numerical integration using the non-thermal distribution $f$ obtained in our computation. Using the entropy densities of the thermal and non-thermal states one can derive the thermalization time $t_{\text{th}}=0.8/\lambda_{L}(T)$, by integrating over Eq. 15. Thus starting from a highly occupied gluon state in a non-thermal scaling regime, a thermal state at $T\simeq 600$ MeV can be reached within a time, $t_{\text{th}}\sim 0.50(3)$ fm/c, whereas it would take $t_{\text{th}}\sim 0.70(5)$ fm/c to reach a thermal state at $T\simeq 450$ MeV.
6 Outlook
In this letter, we provide a detailed understanding of the dynamical properties of soft (magnetic) modes of non-Abelian SU(2) gauge theory as a function of temperature as well as in a particular non-thermal attractor state. These magnetic modes interact non-perturbatively even at asymptotically high temperatures [45] and influence dynamical properties like the sphaleron transition rate [28, 46], and other transport properties of the non-Abelian plasma [47]. We show here that these modes play an important role in the process of thermalization in gauge theories as well. In a high temperature SU(2) plasma, where a clean separation between the hard (ultraviolet momentum) modes and these soft modes is possible, resulting in an effective stochastic description of the latter [28], we demonstrate the chaotic behavior inherent in the system by extracting the spectra of positive Lyapunov exponents. This allows us to measure the Kolmogorov-Sinai entropy rate which determines how fast the entropy flows into the system eventually attaining a maximum. We have used this rate to measure a thermalization time $\sim 0.7$-$0.5$ fm/c to reach to a thermal state at $450$-$600$ MeV starting from a particular non-thermal state of soft gluons which shows a particular self-similar behavior [14]. The only assumption that goes into our calculation is the conservation of energy density of the soft modes during the entire evolution of the system. Interestingly our estimate of the thermalization time in this simple non-expanding system of overoccupied soft gluon modes is consistent with the early switch-on time for hydrodynamics necessary to describe flow observables in a heavy-ion collision event [48] which is described by a glasma initial condition in an expanding spacetime geometry. However our result for the thermalization time is smaller than typical estimates ($\gtrsim 2.5$ fm/c) based on perturbative scattering process [49].
In our approach we can only predict the time taken by the soft modes to thermalize if the initial non-thermal and the final thermal states are known. However understanding the microscopic mechanism that drives the system away from the self-similar non-thermal scaling regime would require a more detailed understanding of how hard particles are created in this regime and how their interactions with the soft modes can be formulated within an effective theory. Possible mechanisms include the onset of plasma instabilities  [50, 51, 52, 53, 54, 55, 13, 14, 56], and the process of thermalization is an inherently non-perturbative phenomenon [57]. We would like to address these topics in a future work.
7 Acknowledgements
We are grateful to Sumilan Banerjee, Rajiv Gavai, Sören Schlichting and Ravi Shanker for discussions related to this work. We also thank the anonymous Referee for several suggested improvements. The authors acknowledge the computing time provided by the High Performance Computing Center at the Institute of Mathematical Sciences.